A lab has a 20% acid solution and a 50% acid solution. How many liters of each are required to obtain 600 liters of a 30% acid solution?
Answer by nerdybill(7384)

A lab has a 20% acid solution and a 50% acid solution. How many liters of each are required to obtain 600 liters of a 30% acid solution?
.
Let x = liters of 20% acid solution
then
600-x = liters of 50% acid solution
.
.20x + .50(600-x) = .30(600)
.20x + 300 - .50x = 180
300 - .30x = 180
300 = .30x + 180
120 = .30x
400 liters = x (amount of 20% acid solution)
.
Amount of 50% acid solution:
600-x = 600-400 = 200 liters